# Reactor-based Search for Axion-Like Particles using CsI(Tl) Detector

### v2

@@ Abstract
[inst4]organization=DEPARTMENT OF PHYSICS & ASTRONOMY, Northwestern University,addressline=2145 Sheridan Road, Tech F165 Evanston, city=Illinois, postcode=60208-3112, country=USA
@@ 1 Introduction
Various astrophysical observations based on the behavior of galaxies in the galaxy clusters [1, 2], rotational velocity curves of stars in galaxies [3], gravitational lensing [4] etc.. have confirmed the existence of non-luminous matter in the universe or the so-called dark matter. About 83% of the matter content of the universe is in the form of dark matter. Among Dark Matter Candidates, Axions have dual motivations, besides being one of the candidates for dark matter in the Universe, they also solve the strong CP problem of the Standard Model.
@@
In this experiment, we focus on the reactor-based search for Axions/Axion-Like Particles (ALPs). These investigations have the ability to explore MeV mass range ALPs by using the large number of photons generated in the nuclear reactor core. Once the axions/ALPs are produced at the reactor core [31], if they travel unimpeded to the detection site, they can scatter off the detector material via inverse Primakoff or inverse Compton-like scatterings to produce photons or decay into di-photons or electron-positron pairs which are then detected using CsI(Tl) scintillator detectors providing strong constraints on ALP-photon coupling.

### v3

@@ 1 Introduction
Various astrophysical observations based on the behavior of galaxies in the galaxy clusters , rotational velocity curves of stars in galaxies, gravitational lensing etc.. have confirmed the existence of non-luminous matter in the universe or the so-called dark matter which among Dark Matter Candidates, Axions have dual motivations, besides being one of the candidates for dark matter, they also solve the strong CP problem of the Standard Model [1, 2, 3, 4, 5].
@@
In this experiment, we focus on the reactor-based search for Axions/Axion-Like Particles (ALPs). Unlike traditional methods which primarily focus on axion-photon coupling, this method enable us to investigate the potential of detecting ALPs through their production in gamma-induced scattering processes and their subsequent interactions with detector materials which leverage low-threshold detectors at reactor-based neutrino experiments to enhance the sensitivity to ALP-photon and ALP-electron couplings. These investigations have the ability to explore MeV mass range ALPs by using the large number of photons generated in the nuclear reactor core. Once the axions/ALPs are produced at the reactor core [1], if they travel unimpeded to the detection site, they can scatter off the detector material via inverse Primakoff or inverse Compton-like scatterings to produce photons or decay into di-photons or electron-positron pairs which are then detected using CsI(Tl) scintillator detectors providing strong constraints on ALP-photon coupling.